Derive the equations of motion for a uniformly accelerated body.



1
Expert's answer
2021-02-25T11:28:22-0500
"a=\frac{v-v_0}{t}\\v=v_0+at\\s=\int vdt=\int (v_0+at)dt\\s=s_0+vt+\frac{1}{2}at^2"